# The DREAMS Project: Disentangling the Impact of Halo-to-Halo Variance and Baryonic Feedback on Milky Way Dark Matter Density Profiles

### v1

@@ 3.2.1 Scale Density
We now examine how the best-fit gNFW parameters change as a function of the DREAMS astrophysics and cosmology variations. Figure 3 shows the scale density $\rho_{s}$ as a function of the five parameters, as well as the halo mass. More specifically, we show the predictions made by an emulator trained to learn the one-parameter dependence of $\rho_{s}$ (see Section 2.3 for details on the training and hyperparameter optimization). The advantage of this approach, compared to showing all the fits to the profiles in Figure 2, is that we can directly parameterize the dependence of $\rho_{s}$ on a parameter-by-parameter basis, instead of sampling from realizations that have multi-parameter dependencies. We therefore keep the parameters fixed to their fiducial TNG value and halo mass fixed to $10^{12}~{\rm M}_{\odot}$, except when explicitly varied. The solid lines are the average predictions of the ensemble of ten emulators. Each shaded region represents the $1\sigma$ uncertainty of the ensemble of emulators via Equation 12, which gives a proxy for the intrinsic halo-to-halo scatter at a given parameter point. Finally, the vertical solid red line at the top of each panel marks the fiducial TNG parameter (from Table 1) and the default halo mass.
@@
We find that $\rho_{s}$ has some dependency on all five of the simulation parameters and halo mass, albeit to a different extent in each case. The parameter that holds the most importance in setting the scale density is $\kappa_{w}$ (top-middle panel). In fact, $\log\rho_{s}$ is roughly inversely proportional to $\log\kappa_{w}$, ranging from values of $\rho_{s}=10^{7.2}~{\rm M}_{\odot}\,{\rm kpc}^{-3}$ at the lowest $\kappa_{w}$ values to $\rho_{s}=10^{6.8}~{\rm M}_{\odot}\,{\rm kpc}^{-3}$ at the highest values of $\kappa_{w}$, a change of $\sim 0.4$ dex. In comparison, the average halo-to-halo variance for $\kappa_{w}$ is also $\sim 0.4$ dex (shaded region). Thus, the impact of $\kappa_{w}$ is comparable to that of the intrinsic scatter of the sample. Beyond $\kappa_{w}$, none of the other simulation parameters drive variations comparable to the average halo-to-halo variations. The total variation in the mean relation for $\bar{e}_{w}$ is $\sim 0.1$ dex, $\epsilon_{f,\,\mathrm{high}}$ is $\sim 0.2$ dex (most of which is at the high AGN feedback variations, see discussion in Section 3.3), $\Omega_{\rm M}$ is $\sim 0.1$ dex, and $\sigma_{8}$ is $\sim 0.15$ dex. Compared to the typical halo-to-halo variation of $\sim 0.2$ dex, we conclude that none of these parameters is crucial in setting the normalization of the dark matter density profiles. Similarly, despite there being a somewhat coherent inverse relationship between halo mass and $\log(\rho_{s})$ (bottom-right panel of Figure 3), halo mass does not drive variations more significant than halo-to-halo variation over the range of our Milky Way-mass halos.
@@
3.2.2 Profile Shape Parameters
@@
When increasing $\bar{e}_{w}$, all the shape parameters tend towards the canonical NFW values. Recall that the NFW profile is the empirical prediction from DMO simulations (Navarro, Frenk, & White, 1997). What is happening here is that these extremely high wind energy variations of the TNG model prevent the growth of any significant stellar component (which we show explicitly in Section 3.3), thus the halos are less dominated by the presence of baryons and behave more like collisionless DMO simulations. It should be noted, though, that the gNFW shape parameters only tend towards the NFW values and do not actually reach them. Thus, while the stronger feedback is preventing the growth of the stellar mass, it is not entirely prevented. We explore the stellar mass growth (or lack thereof) further in the next section and in Rose et al. (In Preparation).
@@
3.3 Central Dark Matter Mass Growth
@@
defined as the logarithmic ratio of dark matter mass enclosed at $0.01R_{200}$ from hydro-to-DMO (similar to the metric used in Rose et al. 2023). This quantity provides a metric for the changes in the inner dark matter distribution relative to a DMO baseline. Dark matter will contract because of the presence of the baryons (see also discussion in Section 4.2), and thus $\Gamma_{0.01}$ should be sensitive to the detailed behavior set by the feedback variations.
Figure 5 shows the central mass growth as a function of the three DREAMS baryonic feedback parameters. As in Section 3.2, we train an ensemble of emulators to learn the single parameter dependency of the central mass growth (see Section 2.3 for more details on the emulator training process) and report the average (solid line) and halo-to-halo variation (shaded region) for $M_{\rm Halo}=10^{12}~{\rm M}_{\odot}$. All three astrophysics parameters impact the central mass growth of the halos. The total halo-to-halo variation is $\sim 0.2$ dex (i.e., $\pm 0.1$ dex standard deviation), whereas we find variations of $\sim 0.5$ dex, $\sim 0.4$ dex, and $\sim 0.2$ dex for $\bar{e}_{w}$, $\kappa_{w}$, and $\epsilon_{f,\,\mathrm{high}}$, respectively. Conversely, there are no strong trends with either $\Omega_{\rm M}$ or $\sigma_{8}$ ($\lesssim 0.05$ dex) and thus do not report them in Figure 5. Similarly, variations to the halo mass have negligible impact on the results; Figure 5 also shows the average trend for $M_{\rm Halo}=10^{11.8}~{\rm M}_{\odot}$ (dashed line) and $10^{12.2}~{\rm M}_{\odot}$ (dotted line), noting that these mass ranges have comparable halo-to-halo variance as the $10^{12.0}~{\rm M}_{\odot}$ range.
@@
The trends in mass growth with supernova physics are qualitatively similar in different halo mass hosts, with normalization shifts associated with evolution in the stellar mass–halo mass relation (shown in Figure 6 as a dashed line for $M_{\rm Halo}=10^{11.8}~{\rm M}_{\odot}$, solid for $10^{12.0}~{\rm M}_{\odot}$, and dotted for $10^{12.2}~{\rm M}_{\odot}$; see also Rose et al. 2025b). In more detail, the variation in stellar mass is slightly larger for less massive hosts: e.g., >1>1 dex in $10^{11.8}~{\rm M}_{\odot}$ halos and $\sim 0.8$ dex in $10^{12.2}~{\rm M}_{\odot}$ halos for $\bar{e}_{w}$ variations, with similar trends, albeit to a lesser extent, with $\kappa_{w}$. Thus, both the energy and wind speed variations are less effective at reducing the stellar mass of the central galaxy in more massive hosts, likely because ejecting material out of a deeper potential via supernova feedback is more difficult. Regardless, the difference in stellar mass growth in the low- versus high-mass halos for supernova energy variations ($\bar{e}_{w}$) likely also explains the (very minor) differences seen in the $\Gamma_{0.01}$ ratios at high $\bar{e}_{w}$ arising from variations in halo mass.
@@
Thus, higher values of $\epsilon_{f,\,\mathrm{high}}$ feedback would not output more total thermal energy over their lifetimes than lower values (shown explicitly in Appendix C, Figure C1). This limiting efficacy of AGN would also be consistent with the (slight) density enhancement seen in the inner regions in the top-right panel of Figure 2.
4 Discussion
4.1 Dark Matter-Only Simulations
Figure 7: Density Profiles from Dark Matter-Only Simulations. The left-hand panel is the same as Figure 2, but for the Dark Matter-Only (DMO) simulations. It shows the DMO density profiles for each simulation normalized by $R_{200}$. The green solid line and shaded region is the weighted mean and standard deviation (see Section 2.4) from the DMO simulations, while the dashed pink line and shaded region is the weighted mean and standard deviation from the hydro simulations (i.e., the same as in Figure 2). The right-hand panels all show the ratio of the hydrodynamic simulations’ density to the corresponding DMO simulation for the same parameter ranges as the right-hand panels of Figure 2. We find that supernova feedback variations drive the deviations (or lack thereof) from hydrodynamic to DMO simulations.
As mentioned in Section 2.1, each hydrodynamic simulation in DREAMS has a corresponding DMO counterpart, which has the equivalent initial conditions. We can therefore make systematic comparisons on the impact of baryons as a function of the varied parameters in the simulations.
The left-hand panel of Figure 7 shows the density profiles from the DMO DREAMS simulations. Each thin gray line is an individual halo, while the solid green line and shaded region are the mean and standard deviation of the individual profiles. For comparison purposes, each DMO halo is assigned the same weight as its hydro counterpart, despite the baryonic feedback parameters not impacting the DMO simulations. As a point of reference, we also overplot the densities for the hydro simulations rom Figure 2. Both the mean and spread of the DMO density profiles are generally less than that of the hydro simulations in the inner radii but are roughly consistent at $\gtrsim 0.1R_{200}$. The spread at $0.01R_{200}$ is $0.22$ dex, which decreases to $0.13$ dex at $0.1R_{200}$ before increasing back to $0.34$ dex at $R_{200}$ due to the presence of subhalos. The larger spread at small radii cements the notion that the feedback variations are the dominant drivers of density profile variations in these inner regions. Moreover, the agreement at larger radii ($0.1R_{200}$) further suggests that intrinsic halo-to-halo variations drive the scatter in this regime.
@@ 4.1 Dark Matter-Only Simulations
In general, the density in the central region of the halo is significantly smaller for the DMO simulations relative to their hydro counterparts. The ratio of $\rho_{\rm hydro}$ to $\rho_{\rm DMO}$ at small radii ($\lesssim 0.02R_{200}$) indicates a steeper inner profile in the hydro simulations—in agreement with the results of Section 3.2 and from the EAGLE simulations (Schaller et al., 2015). At larger radii ($\gtrsim 0.1R_{200}$), the hydro and DMO profiles appear roughly consistent. The degree to which the density increases in the inner region is sensitive to the supernova feedback parameters. In particular, the supernova wind energy ($\bar{e}_{w}$) has a very strong impact relative to the DMO simulations. The lower range of $\bar{e}_{w}$ values leads to a factor of $\sim 3$44 increase in the density compared to DMO, whereas the results for the upper range are roughly consistent with the DMO simulations. This makes sense because the high-energy supernova should prevent baryon content from forming, leading to similar results as the DMO simulations. The supernova wind speed ($\kappa_{w}$) shows a similar trend, albeit to a smaller extent, of increased feedback driving the density profiles closer to the DMO simulations. The upper $10\%$ of $\log(\kappa_{w})$ increases the central density by only a factor of $\sim 1.5$ on average, whereas the highest $10\%$ change the density by a factor of >4>4. On the other hand, the AGN feedback and cosmology variations do not have any significant impact relative to the DMO simulations.
We note that the ratio of $\rho_{\rm hydro}/\rho_{\rm DMO}$ only rarely dips below unity in the inner radii of any of our halos, which contrasts results from the FIRE-2 simulations by McKeown et al. (2022), which show that the addition of baryon physics can suppress the inner density by factors of $\sim 2$. The difference between our work and McKeown et al. (2022) likely stems from the implementation of feedback in FIRE-2. One key difference in the FIRE model is the existence of strong, time-variable stellar feedback-driven winds (Muratov et al., 2015), which can perturb the galactic potential and “core” density profiles (Governato et al., 2010; Pontzen & Governato, 2012; Mostow et al., 2024). We discuss the role of bursty feedback more in Section 4.2.1.
4.2 Contraction of Dark Matter Halos
@@
is the orbit-averaged radius for elliptical orbits, $A$ is a normalization constant, and $w$ is a slope (in log space) parameter that together control how the instantaneous orbital radius $r$ is mapped onto its orbit-averaged value. These fit parameters are determined empirically on a halo-by-halo basis (see discussion in Appendix D).
@@
where the ‘DMO’/‘Hydro’ superscript corresponds to the particular simulation that the enclosed mass pertains to. [^7] Recall from Section 2.1 that the DMO simulations include an $\Omega_{b}$ fraction of baryons, which are treated as collisionless particles. This additional collisionless mass effectively increases the total dark matter mass in the DMO simulations by $\Omega_{b}$. To account for this “extra” mass, we make the following correction: $M_{\rm DM}^{\rm DMO,\,corrected}=M_{\rm DM}^{\rm DMO}\times\left(\frac{\Omega_{\rm M}-\Omega_{b}}{\Omega_{\rm M}}\right)~\,$, (21) where $M_{\rm DM}^{\rm DMO}$ is the mass-enclosed profile for the DMO simulation. From this point onwards, we refer to the “corrected” quantity when referencing $M_{\rm DM}^{\rm DMO}$. Here, the total enclosed mass of the DMO halo is rescaled by
@@
Figure 8: Ratio of Estimated Mass Contraction to Hydrodynamic Simulation Mass Profiles. The ratio of the predicted adiabatically contracted dark matter mass profile ($M_{\rm DM}^{\rm AC}$; using the methods outlined in Gnedin et al. 2004) to the dark matter mass profile of the full hydrodynamic simulations ($M_{\rm DM}^{\rm Hydro}$), shown as a function of radius (rescaled by $R_{200}$. Individual galaxies are shown by the gray lines. The solid line represents the weighted mean of the distribution, while the shaded region is the weighted spread (see Section 2.4). The right-hand panels show the ratio for the same parameter ranges as the right-hand panels of Figure 2. The level to which the halos agree with the adiabatically contracting halo model is mostly independent of astrophysics or cosmology parameter variations at radii $\gtrsim 0.03R_{200}$ (with the exception of $\bar{e}_{w}$).
@@
A small fraction of halos ($27$) deviate significantly ($\gg 10\%$) from the adiabatic contraction model at $R_{200}$. The large deviation at this radius is surprising considering that part of the calculation above is to renormalize the DMO profile based on masses at the virial radius (e.g., Equation 22). Upon further inspection, all of the cases where the profiles deviate significantly have a merger/large accretion event crossing the virial radius at $z\sim 0$. The exact timing of this event is different between the DMO and hydrodynamic simulations (which could be for a number of reasons, physical or numerical; Genel et al. 2019, Pakmor et al. 2025). The required renormalization of the DMO mass profile can therefore be systematically over- or under-estimated in situations where the infalling material crosses the virial radius at different times, leading to a failure of the adiabatically contracting halo model. Indeed, a halo undergoing a merger/large accretion event is likely to not be adiabatically (or even quasi-adiabatically) contracting (Velmani & Paranjape, 2023).
4.2.1 Comparison with Bursty Feedback Model

### v2

@@ III.2.1 Scale Density
We now examine how the best-fit gNFW parameters change as a function of the DREAMS astrophysics and cosmology variations. Figure 3 shows the scale density $\rho_{s}$ as a function of the five parameters, as well as the halo mass. More specifically, we show the predictions made by an emulator trained to learn the one-parameter dependence of $\rho_{s}$ (see Section II.3 for details on the training and hyperparameter optimization). The advantage of this approach, compared to showing all the fits to the profiles in Figure 2, is that we can directly parameterize the dependence of $\rho_{s}$ on a parameter-by-parameter basis, instead of sampling from realizations that have multi-parameter dependencies. We therefore keep the parameters fixed to their fiducial TNG value and halo mass fixed to $10^{12}~{\rm M}_{\odot}$, except when explicitly varied. The solid lines are the average predictions of the ensemble of ten emulators. Each shaded region represents the $1\sigma$ uncertainty of the ensemble of emulators via Equation 12, which gives a proxy for the intrinsic halo-to-halo scatter at a given parameter point. Finally, the vertical solid red line at the top of each panel marks the fiducial TNG parameter (from Table 1) and the default halo mass.
@@
We find that $\rho_{s}$ has some dependency on all five of the simulation parameters and halo mass, albeit to a different extent in each case. The parameter that holds the most importance in setting the scale density is $\kappa_{w}$ (top-middle panel). In fact, $\log\rho_{s}$ is roughly inversely proportional to $\log\kappa_{w}$, ranging from values of $\rho_{s}=10^{7.2}~{\rm M}_{\odot}\,{\rm kpc}^{-3}$ at the lowest $\kappa_{w}$ values to $\rho_{s}=10^{6.8}~{\rm M}_{\odot}\,{\rm kpc}^{-3}$ at the highest values of $\kappa_{w}$, a change of $\sim 0.4$ dex. In comparison, the average halo-to-halo variance for $\kappa_{w}$ is also $\sim 0.4$ dex (shaded region). Thus, the impact of $\kappa_{w}$ is comparable to that of the intrinsic scatter of the sample. Beyond $\kappa_{w}$, none of the other simulation parameters drive variations comparable to the average halo-to-halo variations. The total variation in the mean relation for $\bar{e}_{w}$ is $\sim 0.1$ dex, $\epsilon_{f,\,\mathrm{high}}$ is $\sim 0.2$ dex (most of which is at the high AGN feedback variations, see discussion in Section III.3), $\Omega_{\rm M}$ is $\sim 0.1$ dex, and $\sigma_{8}$ is $\sim 0.15$ dex. Compared to the typical halo-to-halo variation of $\sim 0.2$ dex, we conclude that none of these parameters is crucial in setting the normalization of the dark matter density profiles. Similarly, despite there being a somewhat coherent inverse relationship between halo mass and $\log(\rho_{s})$ (bottom-right panel of Figure 3), halo mass does not drive variations more significant than halo-to-halo variation over the range of our Milky Way-mass halos.
@@
III.2.2 Profile Shape Parameters
@@
When increasing $\bar{e}_{w}$, all the shape parameters tend towards the canonical NFW values. Recall that the NFW profile is the empirical prediction from DMO simulations (Navarro, Frenk, & White, 1997). What is happening here is that these extremely high wind energy variations of the TNG model prevent the growth of any significant stellar component (which we show explicitly in Section III.3), thus the halos are less dominated by the presence of baryons and behave more like collisionless DMO simulations. It should be noted, though, that the gNFW shape parameters only tend towards the NFW values and do not actually reach them. Thus, while the stronger feedback is preventing the growth of the stellar mass, it is not entirely prevented. We explore the stellar mass growth (or lack thereof) further in the next section and in Rose et al. (2025c).
@@
III.3 Central Dark Matter Mass Growth
@@
defined as the logarithmic ratio of dark matter mass enclosed at $0.01R_{200}$ from hydro-to-DMO (similar to the metric used in Rose et al. 2023). This quantity provides a metric for the changes in the inner dark matter distribution relative to a DMO baseline. Dark matter will contract because of the presence of the baryons (see also discussion in Section IV.2), and thus $\Gamma_{0.01}$ should be sensitive to the detailed behavior set by the feedback variations.
Figure 5 shows the central mass growth as a function of the three DREAMS baryonic feedback parameters. As in Section III.2, we train an ensemble of emulators to learn the single parameter dependency of the central mass growth (see Section II.3 for more details on the emulator training process) and report the average (solid line) and halo-to-halo variation (shaded region) for $M_{\rm Halo}=10^{12}~{\rm M}_{\odot}$. All three astrophysics parameters impact the central mass growth of the halos. The total halo-to-halo variation is $\sim 0.2$ dex (i.e., $\pm 0.1$ dex standard deviation), whereas we find variations of $\sim 0.5$ dex, $\sim 0.4$ dex, and $\sim 0.2$ dex for $\bar{e}_{w}$, $\kappa_{w}$, and $\epsilon_{f,\,\mathrm{high}}$, respectively. Conversely, there are no strong trends with either $\Omega_{\rm M}$ or $\sigma_{8}$ ($\lesssim 0.05$ dex) and thus do not report them in Figure 5. Similarly, variations to the halo mass have negligible impact on the results; Figure 5 also shows the average trend for $M_{\rm Halo}=10^{11.8}~{\rm M}_{\odot}$ (dashed line) and $10^{12.2}~{\rm M}_{\odot}$ (dotted line), noting that these mass ranges have comparable halo-to-halo variance as the $10^{12.0}~{\rm M}_{\odot}$ range.
@@
The trends in mass growth with supernova physics are qualitatively similar in different halo mass hosts, with normalization shifts associated with evolution in the stellar mass–halo mass relation (shown in Figure 6 as a dashed line for $M_{\rm Halo}=10^{11.8}~{\rm M}_{\odot}$, solid for $10^{12.0}~{\rm M}_{\odot}$, and dotted for $10^{12.2}~{\rm M}_{\odot}$; see also R25b). In more detail, the variation in stellar mass is slightly larger for less massive hosts: e.g., >1>1 dex in $10^{11.8}~{\rm M}_{\odot}$ halos and $\sim 0.8$ dex in $10^{12.2}~{\rm M}_{\odot}$ halos for $\bar{e}_{w}$ variations, with similar trends, albeit to a lesser extent, with $\kappa_{w}$. Thus, both the energy and wind speed variations are less effective at reducing the stellar mass of the central galaxy in more massive hosts, likely because ejecting material out of a deeper potential via supernova feedback is more difficult. Regardless, the difference in stellar mass growth in the low- versus high-mass halos for supernova energy variations ($\bar{e}_{w}$) likely also explains the (very minor) differences seen in the $\Gamma_{0.01}$ ratios at high $\bar{e}_{w}$ arising from variations in halo mass.
@@
Thus, higher values of $\epsilon_{f,\,\mathrm{high}}$ feedback would not output more total thermal energy over their lifetimes than lower values (shown explicitly in Appendix B, Figure C1). This limiting efficacy of AGN would also be consistent with the (slight) density enhancement seen in the inner regions in the top-right panel of Figure 2.
IV Discussion
IV.1 Dark Matter-Only Simulations
Figure 7: Density Profiles from Dark Matter-Only Simulations. The left-hand panel is the same as Figure 2, but for the Dark Matter-Only (DMO) simulations. It shows the DMO density profiles for each simulation normalized by $R_{200}$. The green solid line and shaded region is the weighted mean and standard deviation (see Section II.4) from the DMO simulations, while the dashed pink line and shaded region is the weighted mean and standard deviation from the hydro simulations (i.e., the same as in Figure 2). The right-hand panels all show the ratio of the hydrodynamic simulations’ density to the corresponding DMO simulation for the same parameter ranges as the right-hand panels of Figure 2. We find that supernova feedback variations drive the deviations (or lack thereof) from hydrodynamic to DMO simulations.
As mentioned in Section II.1, each hydrodynamic simulation in DREAMS has a corresponding DMO counterpart, which has the equivalent initial conditions. We can therefore make systematic comparisons on the impact of baryons as a function of the varied parameters in the simulations.
The left-hand panel of Figure 7 shows the density profiles from the DMO DREAMS simulations. Each thin gray line is an individual halo, while the solid green line and shaded region are the mean and standard deviation of the individual profiles. For comparison purposes, each DMO halo is assigned the same weight as its hydro counterpart, despite the baryonic feedback parameters not impacting the DMO simulations. As a point of reference, we also overplot the densities for the hydro simulations rom Figure 2. Both the mean and spread of the DMO density profiles are generally less than that of the hydro simulations in the inner radii but are roughly consistent at $\gtrsim 0.1R_{200}$. The spread at $0.01R_{200}$ is $0.22$ dex, which decreases to $0.13$ dex at $0.1R_{200}$ before increasing back to $0.34$ dex at $R_{200}$ due to the presence of subhalos. The larger spread at small radii cements the notion that the feedback variations are the dominant drivers of density profile variations in these inner regions. Moreover, the agreement at larger radii ($0.1R_{200}$) further suggests that intrinsic halo-to-halo variations drive the scatter in this regime. 666 We note that we normalize the radial profiles of both the DMO and hydrodynamic simulations by $R_{200}$. In detail, $R_{200}$ is not the same radius in the pairs of simulations; however, we find that our results are qualitatively unimpacted by this difference. We find similar results when normalizing both profiles by their respective best-fit $\rho_{0}$ and $r_{s}$ gNFW parameters.
@@ IV.1 Dark Matter-Only Simulations
In general, the density in the central region of the halo is significantly smaller for the DMO simulations relative to their hydro counterparts. The ratio of $\rho_{\rm hydro}$ to $\rho_{\rm DMO}$ at small radii ($\lesssim 0.02R_{200}$) indicates a steeper inner profile in the hydro simulations—in agreement with the results of Section III.2 and from the EAGLE simulations (Schaller et al., 2015). At larger radii ($\gtrsim 0.1R_{200}$), the hydro and DMO profiles appear roughly consistent. The degree to which the density increases in the inner region is sensitive to the supernova feedback parameters. In particular, the supernova wind energy ($\bar{e}_{w}$) has a very strong impact relative to the DMO simulations. The lower range of $\bar{e}_{w}$ values leads to a factor of $\sim 3$44 increase in the density compared to DMO, whereas the results for the upper range are roughly consistent with the DMO simulations. This makes sense because the high-energy supernova should prevent baryon content from forming, leading to similar results as the DMO simulations. The supernova wind speed ($\kappa_{w}$) shows a similar trend, albeit to a smaller extent, of increased feedback driving the density profiles closer to the DMO simulations. The upper $10\%$ of $\log(\kappa_{w})$ increases the central density by only a factor of $\sim 1.5$ on average, whereas the highest $10\%$ change the density by a factor of >4>4. On the other hand, the AGN feedback and cosmology variations do not have any significant impact relative to the DMO simulations.
We note that the ratio of $\rho_{\rm hydro}/\rho_{\rm DMO}$ only rarely dips below unity in the inner radii of any of our halos, which contrasts results from the FIRE-2 simulations by McKeown et al. (2022), which show that the addition of baryon physics can suppress the inner density by factors of $\sim 2$. The difference between our work and McKeown et al. (2022) likely stems from the implementation of feedback in FIRE-2. One key difference in the FIRE model is the existence of strong, time-variable stellar feedback-driven winds (Muratov et al., 2015), which can perturb the galactic potential and “core” density profiles (Governato et al., 2010; Pontzen & Governato, 2012; Mostow et al., 2024). We discuss the role of bursty feedback more in Section IV.2.1.
IV.2 Contraction of Dark Matter Halos
@@
is the orbit-averaged radius for elliptical orbits, $A$ is a normalization constant, and $w$ is a slope (in log space) parameter that together control how the instantaneous orbital radius $r$ is mapped onto its orbit-averaged value. These fit parameters are determined empirically on a halo-by-halo basis (see discussion in Appendix C).
@@
where the ‘DMO’/‘Hydro’ superscript corresponds to the particular simulation that the enclosed mass pertains to. [^8] Recall from Section II.1 that the DMO simulations include an $\Omega_{b}$ fraction of baryons, which are treated as collisionless particles. This additional collisionless mass effectively increases the total dark matter mass in the DMO simulations by $\Omega_{b}$. To account for this “extra” mass, we make the following correction: $M_{\rm DM}^{\rm DMO,\,corrected}=M_{\rm DM}^{\rm DMO}\times\left(\frac{\Omega_{\rm M}-\Omega_{b}}{\Omega_{\rm M}}\right)~\,$, (21) where $M_{\rm DM}^{\rm DMO}$ is the mass-enclosed profile for the DMO simulation. From this point onwards, we refer to the “corrected” quantity when referencing $M_{\rm DM}^{\rm DMO}$. Here, the total enclosed mass of the DMO halo is rescaled by
@@
Figure 8: Ratio of Estimated Mass Contraction to Hydrodynamic Simulation Mass Profiles. The ratio of the predicted adiabatically contracted dark matter mass profile ($M_{\rm DM}^{\rm AC}$; using the methods outlined in Gnedin et al. 2004) to the dark matter mass profile of the full hydrodynamic simulations ($M_{\rm DM}^{\rm Hydro}$), shown as a function of radius (rescaled by $R_{200}$. Individual galaxies are shown by the gray lines. The solid line represents the weighted mean of the distribution, while the shaded region is the weighted spread (see Section II.4). The right-hand panels show the ratio for the same parameter ranges as the right-hand panels of Figure 2. The level to which the halos agree with the adiabatically contracting halo model is mostly independent of astrophysics or cosmology parameter variations at radii $\gtrsim 0.03R_{200}$ (with the exception of $\bar{e}_{w}$).
@@
A small fraction of halos ($27$ out of the 1024; $2.6\%$) deviate significantly ($\gg 10\%$) from the adiabatic contraction model at $R_{200}$. The large deviation at this radius is surprising considering that part of the calculation above is to renormalize the DMO profile based on masses at the virial radius (e.g., Equation 22). Upon further inspection, all of the cases where the profiles deviate significantly have a merger/large accretion event crossing the virial radius at $z\sim 0$. The exact timing of this event is different between the DMO and hydrodynamic simulations (which could be for a number of reasons, physical or numerical; Genel et al. 2019, Pakmor et al. 2025). The required renormalization of the DMO mass profile can therefore be systematically over- or under-estimated in situations where the infalling material crosses the virial radius at different times, leading to a failure of the adiabatically contracting halo model. Indeed, a halo undergoing a merger/large accretion event is likely to not be adiabatically (or even quasi-adiabatically) contracting (Velmani & Paranjape, 2023).
IV.2.1 Comparison with Bursty Feedback Model